# Real-world models for multiple term structures: a unifying HJM semimartingale framework

### v2

@@ 4.2. Well-posedness of the real-world HJMM SPDE
As a consequence of Theorem D.1 and Remark 4.9, the space $(H,\|\cdot\|_{H})$ is a separable Hilbert space and the shift semigroup $(S_{t})_{t\geq 0}$ is a $C_{0}$-semigroup on $H$ with generator $A$ given by $Ah=h^{\prime}$, for all $h\in D(A)$. Moreover, [5, Lemma 3.5] implies that the shift semigroup $(S_{t})_{t\geq 0}$ is pseudo-contractive. By Proposition 4.12, conditions (4.13)–(4.18) are satisfied. The claim then follows by Theorem 4.2. ∎
@@ 4.3. Monotonicity of term structures
where we recall that the state space is given by $H=H_{\rho}^{m+1}$. We say that the closed convex cone $K$ is invariant for the SPDE (4.25) if, for each $h_{0}\in K$, we have $\eta\in K$ up to an evanescent set, where $\eta$ denotes the mild solution to the SPDE (4.25) with $\eta_{0}=h_{0}$. The invariance of closed convex cones for SPDEs has been investigated in a general setup in [51, 52] and in [25] for the special case where $K$ is the cone of non-negative functions. The proof of the following theorem is analogous to that of [52, Corollary 8.6] and, therefore, omitted.
@@
[…]
@@
for $(\mathbb{P}\otimes{\rm d}t)$-a.e. $(\omega,t)\in\Omega\times\mathbb{R}_{+}$. Recalling that $\psi$ is $(-1,+\infty)$-valued and $c^{i}$ is $[-1,+\infty)$-valued, conditions (4.53), (4.54), (4.56), (4.57) imply that (4.51) is fulfilled. Consequently, in view of Theorem 4.14, the closed convex cone $K$ is invariant for the HJMM SPDE (4.25). ∎
@@ 4.4. Existence of affine realizations
where $L$ is an $\mathbb{R}^{d}$-valued Lévy process, $A$ denotes the generator of a pseudo-contractive $C_{0}$-semigroup $(S_{t})_{t\geq 0}$ on $H$ and $\alpha:\Omega\times\mathbb{R}_{+}\times H\to H$ and $\beta:\Omega\times\mathbb{R}_{+}\times H\to L_{2}(\mathbb{R}^{d},H)$. For each $j=1,\ldots,d$, we introduce the mapping $\beta^{j}:\Omega\times\mathbb{R}_{+}\times H\to H$ given by $\beta^{j}(\omega,t,h):=\beta(\omega,t,h)e_{j}$. We assume that conditions (4.13), (4.14) and (4.16), (4.17) from Theorem 4.2 are satisfied, which ensures existence and uniqueness of mild solutions to the Lévy-driven SPDE (4.59).
@@
Let $V\subset D(A)$ be an $n$-dimensional $A$-invariant subspace such that the following conditions are fulfilled:
@@
$\phi^{x_{0}}(t):=S_{t}x_{0}+\int_{0}^{t}S_{t-s}\alpha_{1}(s){\rm d}s,\quad\text{ for all }t\in\mathbb{R}_{+}.$
@@
see for instance [45, Theorem 4.2.4]. In particular, it follows that $A\phi^{x_{0}}:\mathbb{R}_{+}\to H$ is continuous. Condition (4.62) is satisfied due to (4.66). Furthermore, for all $(\omega,t)\in\Omega\times\mathbb{R}_{+}$ and $v\in V$, we obtain
@@
Suppose that the SPDE (4.59) has an $n$-dimensional affine realization induced by $(\phi,T)$. Let $\ell\in L(H,\mathbb{R}^{n})$ be such that the linear operator $S:=\ell T\in L(\mathbb{R}^{n})$ is an isomorphism and let $R\in L(\mathbb{R}^{n},V)$ be the isomorphism given by $R:=TS^{-1}$. Then, for each $x_{0}\in D(A)$, the following hold:
@@
$Y_{t}^{x_{0}}=S^{-1}\ell(X_{t}^{x_{0}}-\phi^{x_{0}}(t)),\quad\text{ for all }t\in\mathbb{R}_{+}.$
@@
    (4.75) $\varphi^{h_{0},i}(t):=S_{t}h_{0}^{i}-h_{0}^{i}(t)e^{\delta_{i}\cdot}+\frac{c_{i}^{2}}{2\delta_{i}^{2}}\big{(}e^{2\delta_{i}t}-1\big{)}\big{(}e^{\delta_{i}\cdot}-1\big{)}e^{\delta_{i}\cdot},\;\text{ for all }t\in\mathbb{R}_{+}\text{ and }i=0,1,\ldots,m;$
@@
$\phi^{h_{0}}(t)=S_{t}h_{0}+\int_{0}^{t}S_{t-s}\alpha_{1}\,{\rm d}s,\qquad\text{ for all }t\in\mathbb{R}_{+}.$
Note that $S:=\ell T\in L(\mathbb{R}^{m+1})$ is simply given by $S={\rm Id}$. The isomorphism $R:=TS^{-1}\in L(\mathbb{R}^{m+1},V)$ then reduces to $R=T$ and, therefore, $R\ell\in L(H,V)$ is given by
@@
$\int_{0}^{t}S_{t-s}\alpha_{1}\,{\rm d}s=\bigg{(}\frac{c_{i}^{2}}{2\delta_{i}^{2}}\Big{(}\big{(}e^{2\delta_{i}t}-1\big{)}e^{\delta_{i}\cdot}-2\big{(}e^{\delta_{i}t}-1\big{)}\Big{)}e^{\delta_{i}\cdot}\bigg{)}_{i=0,1,\ldots,m}.$
@@
$R\ell\int_{0}^{t}S_{t-s}\alpha_{1}\,{\rm d}s=\bigg{(}\frac{c_{i}^{2}}{2\delta_{i}^{2}}\big{(}e^{2\delta_{i}t}-2e^{\delta_{i}t}+1\big{)}e^{\delta_{i}\cdot}\bigg{)}_{i=0,1,\ldots,m}.$
@@
$\frac{{\rm d}}{{\rm d}\xi}\varphi^{h_{0}}(t)=\bigg{(}S_{t}\frac{{\rm d}}{{\rm d}\xi}h_{0}^{i}-\delta_{i}h_{0}^{i}(t)e^{\delta_{i}\cdot}+\frac{c_{i}^{2}}{2\delta_{i}}\big{(}e^{2\delta_{i}t}-1\big{)}\big{(}2e^{2\delta_{i}\cdot}-e^{\delta_{i}\cdot}\big{)}\bigg{)}_{i=0,1,\ldots,m},$
@@
For each $n\in\mathbb{N}$, let us denote by $\mathcal{X}_{1,[0,n]}:=\{X_{\cdot\wedge n}:X\in\mathcal{X}_{1}\}$ the set of all elements of $\mathcal{X}_{1}$ stopped at $n$. The set $1+\mathcal{X}_{1,[0,n]}$ satisfies the requirements of [40, Definition 1.1], since fork-convexity follows as a consequence of [18, Lemma 2.1] (see also [19, proof of Theorem 4.7]), and is closed in the semimartingale topology by definition. Suppose first that NUPBR holds. Then, for every $n\in\mathbb{N}$, the set $\mathcal{X}_{1}(n)$ is bounded in probability and therefore, by [40, Theorem 1.7], there exists an element $\widehat{X}^{n}\in\mathcal{X}_{1,[0,n]}$ with $\widehat{X}^{n}>-1$ such that $(1+X)/(1+\widehat{X}^{n})$ is a supermartingale, for every $X\in\mathcal{X}_{1,[0,n]}$. For each $n\in\mathbb{N}$, let us denote $Z^{n}:=1/(1+\widehat{X}^{n})$. Similarly as in the proof of [14, Proposition 1], we show that the elements $Z^{n}$, $n\in\mathbb{N}$, can be concatenated into a supermartingale deflator. For all $t\geq 0$, let $n(t):=\min\{n\in\mathbb{N}:n>t\}$ and define the càdlàg process
(A.1) $Z_{t}:=\prod_{k=1}^{n(t)}\frac{Z^{k}_{k\wedge t}}{Z^{k}_{(k-1)\wedge t}},\qquad\text{ for all }t\geq 0.$
@@
Notice that this definition implies that, if $t\in(m-1,m]$ for some $m\in\mathbb{N}$, then $Z_{t}=\prod_{k=1}^{m}(Z^{k}_{k\wedge t}/Z^{k}_{k-1})$. Let $X\in\mathcal{X}_{1}$ and $s<t$. Suppose first that $t\in(s,n(s)]\subseteq(n(s)-1,n(s)]$. In this case, we can compute
@@
$\mathbb{E}\left[Z_{t}(1+X_{t})|\mathcal{F}_{s}\right]$ $=\mathbb{E}\left[\prod_{k=1}^{n(s)}\frac{Z^{k}_{k\wedge t}}{Z^{k}_{k-1}}(1+X_{t})\Bigg{|}\mathcal{F}_{s}\right]$
$=\prod_{k=1}^{n(s)-1}\frac{Z^{k}_{k}}{Z^{k}_{k-1}}\,\mathbb{E}\left[\frac{Z^{n(s)}_{t}}{Z^{n(s)}_{n(s)-1}}(1+X_{t})\Bigg{|}\mathcal{F}_{s}\right]$
$\leq\prod_{k=1}^{n(s)-1}\frac{Z^{k}_{k}}{Z^{k}_{k-1}}\frac{Z^{n(s)}_{s}(1+X_{s})}{{Z^{n(s)}_{n(s)-1}}}=\prod_{k=1}^{n(s)}\frac{Z^{k}_{k\wedge s}}{Z^{k}_{k-1}}(1+X_{s})=Z_{s}(1+X_{s}),$
@@
where the inequality follows from the fact that $Z^{n(s)}(1+X)$ is a supermartingale on $[0,n(s)]$. This shows that the supermartingale property holds for $t\in(s,n(s)]$. Proceeding by induction, to prove that $Z$ is a supermartingale deflator it suffices to show that, if the supermartingale property holds for $t\in(s,n(s)+l]$, for some $l\in\mathbb{N}$, then it also holds for $t\in(n(s)+l,n(s)+l+1]$. To this effect, we compute
$\mathbb{E}\left[Z_{t}(1+X_{t})|\mathcal{F}_{s}\right]$ $=\mathbb{E}\left[\mathbb{E}\left[Z_{t}(1+X_{t})|\mathcal{F}_{n(s)+l}\right]\big{|}\mathcal{F}_{s}\right]$
$=\mathbb{E}\left[Z_{n(s)+l}\,\mathbb{E}\left[\frac{Z^{n(s)+l+1}_{t}}{Z^{n(s)+l+1}_{n(s)+l}}(1+X_{t})\Bigg{|}\mathcal{F}_{n(s)+l}\right]\Bigg{|}\mathcal{F}_{s}\right]$
$\leq\mathbb{E}\left[Z_{n(s)+l}(1+X_{n(s)+l})|\mathcal{F}_{s}\right]\leq Z_{s}(1+X_{s}),$
@@
where the first inequality follows from the supermartingale property of $Z^{n(s)+l+1}(1+X)$ on the interval $[0,n(s)+l+1]$ and the second inequality from the induction step. Since $Z_{0}=1$, this completes the proof that the process $Z$ defined in (A.1) is a supermartingale deflator in the sense of Definition 2.4. We now prove that $1/Z\in 1+\mathcal{X}_{1}$. To this effect, note that
$\widehat{X}_{t}:=\frac{1}{Z_{t}}-1=\prod_{k=1}^{n(t)}\frac{1+\widehat{X}^{k}_{k\wedge t}}{1+\widehat{X}^{k}_{(k-1)\wedge t}}-1,\qquad\text{ for all }t\geq 0.$

### v3

@@ 4.2. Well-posedness of the real-world HJMM SPDE
As a consequence of Theorem D.1 and Remark 4.9, the space $(H,\|\cdot\|_{H})$ is a separable Hilbert space and the shift semigroup $(\mathcal{S}_{t})_{t\geq 0}$ is a $C_{0}$-semigroup on $H$ with generator $A$ given by $Ah=h^{\prime}$, for all $h\in D(A)$. Moreover, [5, Lemma 3.5] implies that the shift semigroup $(\mathcal{S}_{t})_{t\geq 0}$ is pseudo-contractive. By Proposition 4.13, conditions (4.13)–(4.18) are satisfied. The claim then follows by Theorem 4.2. ∎
Example 4.15.
We provide an example of a class of models with unbounded coefficients $\beta$ and $\gamma$ for which Assumption 4.10 is fulfilled. For simplicity, we consider the situation $d=1$ (one-dimensional Wiener process) and $m=0$ (single curve setting) and agree on the notation $H_{\rho}:=H_{\rho}^{1}$, in accordance with Appendix D. First of all, let $\lambda$ and $\psi$ be any processes such that items (i)-(ii) of Assumption 4.10 are satisfied. In particular, we can have $\lambda\equiv 0$ and $\psi\equiv 0$, which corresponds to the situation under a risk-neutral probability $\mathbb{Q}$ (see Section 3.4). In that case, this example yields a class of risk-neutral models not covered by the results of [25], due to the unboundedness of $\beta$ and $\gamma$.
We assume that the diffusive coefficient $\beta:H_{\rho}\to H_{\rho}$ is given by a constant direction volatility
(4.48) $\beta(h)=\Phi_{\beta}(h)\cdot\lambda_{\beta},\quad\text{ for }h\in H_{\rho}$
with $\lambda_{\beta}\in H_{\rho}^{0}$ and a locally Lipschitz functional $\Phi_{\beta}:H_{\rho}\to\mathbb{R}$ satisfying the growth condition
(4.49) $|\Phi_{\beta}(h)|\leq C_{1}\sqrt{1+\|h\|_{\rho}},\quad\text{ for all $h\in H$,}$
for some constant $C_{1}>0$. Then condition (4.30) is satisfied and (4.31) is fulfilled with $M_{\beta}=C_{1}\|\lambda_{\beta}\|_{\rho}$. The functional $\Phi_{\beta}$ can for instance be of the form $\Phi_{\beta}=f\circ\ell$ with a continuous linear functional $\ell\in L(H_{\rho},\mathbb{R})$ and a locally Lipschitz function $f:\mathbb{R}\to\mathbb{R}$ satisfying the growth condition
(4.50) $|f(\xi)|\leq C_{2}\sqrt{1+|\xi|},\quad\text{ for all $\xi\in\mathbb{R}$,}$
for some constant $C_{2}>0$. Indeed, in this case the functional $\Phi_{\beta}$ is locally Lipschitz and the estimate (4.49) is satisfied with $C_{1}=C_{2}\sqrt{1+\|\ell\|}$. Examples for the linear functional $\ell$ are the point evaluations
$\mathcal{J}_{x}:H_{\rho}\to\mathbb{R},\quad\mathcal{J}_{x}(h):=h(x)$
for any $x\in\mathbb{R}_{+}$, and the integral functionals
$\mathcal{I}_{u}:H_{\rho}\to\mathbb{R},\quad\mathcal{I}_{u}(h):=\int_{0}^{u}h(x)\,{\rm d}x.$
for any $u\in\mathbb{R}_{+}$, see [22, Theorem 5.1.1 and Lemma 4.3.1]. Furthermore, for any fixed constant $\eta>0$ an example of a locally Lipschitz function $f:\mathbb{R}\to\mathbb{R}$ satisfying (4.50) is given by
$f(\xi)=\frac{|\xi|}{\sqrt{\eta}}\mathbf{1}_{\{|\xi|\leq\eta\}}+\sqrt{|\xi|}\mathbf{1}_{\{|\xi|>\eta\}},\quad\text{ for }\xi\in\mathbb{R}.$
In order to specify the jump coefficient $\gamma$, we choose another constant $\rho^{\prime}>\rho$. Let $W:\mathbb{R}_{+}\to\mathbb{R}_{+}$ be a Lipschitz continuous function such that
(4.51) $W(r)\leq W_{K_{\rho,\rho^{\prime}}}(r),\quad\text{ for all $r\in\mathbb{R}_{+}$,}$
where we recall that $W_{K_{\rho,\rho^{\prime}}}:\mathbb{R}_{+}\to\mathbb{R}_{+}$ denotes the inverse of the strictly increasing function (4.28) with $K=K_{\rho,\rho^{\prime}}$. We assume that the jump coefficient $\gamma:H_{\rho}\times E\to H_{\rho^{\prime}}$ is given by
$\gamma(h,x)=\Phi_{\gamma}(h,x)\cdot\lambda_{\gamma},\quad\text{ for }(h,x)\in H_{\rho}\times E,$
with $\lambda\in H_{\rho^{\prime}}^{0}$ and a mapping $\Phi_{\gamma}:H_{\rho}\times E\to\mathbb{R}_{+}$ given by
(4.52) $\Phi_{\gamma}(h,x)=c\cdot W(\kappa(x)\Psi(h)),\quad\text{ for }(h,x)\in H_{\rho}\times E,$
where the constant $c\in\mathbb{R}_{+}$ satisfies $c\leq\|\lambda_{\gamma}\|_{\rho^{\prime}}^{-1}$ and the function $\Psi:H_{\rho}\to\mathbb{R}_{+}$ is locally Lipschitz and satisfies the linear growth condition. Then, $\gamma$ is $\mathcal{B}(H_{\rho})\otimes\mathcal{B}(E)$-measurable. Furthermore, using (4.51) and the monotonicity of $W_{K_{\rho,\rho^{\prime}}}$, we have that
$\|\gamma(h,x)\|_{\rho^{\prime}}$ $=\|\lambda_{\gamma}\|_{\rho^{\prime}}\,|\Phi_{\gamma}(h,x)|\leq W(\kappa(x)\Psi(h))\leq W_{K_{\rho,\rho^{\prime}}}(\kappa(x)\Psi(h))\leq W_{K_{\rho,\rho^{\prime}}}\big(C_{\Psi}\kappa(x)(1+\|h\|_{\rho})\big),$
for all $h\in H_{\rho}$ and $x\in E$, where $C_{\Psi}\in\mathbb{R}_{+}$ denotes a growth constant of $\Psi$. This shows that condition (4.33) holds with $N_{\gamma}=C_{\Psi}$. Moreover, for $r\in\mathbb{R}_{+}$ and $h,g\in H_{\rho}$ with $\|h\|_{\rho}\vee\|g\|_{\rho}\leq r$, it holds that
$\|\gamma(h,x)-\gamma(g,x)\|_{\rho^{\prime}}$ $=\|\lambda_{\gamma}\|_{\rho^{\prime}}\,|\Phi_{\gamma}(h,x)-\Phi_{\gamma}(g,x)|\leq|W(\kappa(x)\Psi(h))-W(\kappa(x)\Psi(g))|$
$\leq L_{W}|\kappa(x)\Psi(h)-\kappa(x)\Psi(g)|=L_{W}\kappa(x)|\Psi(h)-\Psi(g)|\leq L_{W}\kappa(x)L_{\Psi}(r)\|h-g\|_{\rho},$
where $L_{W}\in\mathbb{R}_{+}$ denotes a Lipschitz constant of $W$ and $L_{\Psi}$ denotes a Lipschitz function of $\Psi$. This shows the validity of condition (4.32) with $M_{\gamma}(r)=L_{W}\cdot L_{\Psi}(r)$. We have therefore shown that all requirements of Assumption 4.10 are satisfied.
To make the choice of the function $W$ in (4.52) more concrete, let us consider two examples of Lipschitz continuous functions $W:\mathbb{R}_{+}\to\mathbb{R}_{+}$ satisfying (4.51):
  1. (1)
    we can simply take $W:=W_{K_{\rho,\rho^{\prime}}}$, which obviously satisfies condition (4.51). Moreover, by the Leibniz rule, letting $K=K_{\rho,\rho^{\prime}}$, we have that
    $V_{K}^{\prime}(r)=r(1+r)Ke^{Kr}+2re^{Kr}+(1+r)e^{Kr}\geq 1,\quad\text{ for all $r\in\mathbb{R}_{+}$,}$
    where we recall that $V_{K}$ is given by (4.28). Thus, by the inverse mapping theorem for differentiation we have $W^{\prime}(r)\leq 1$ for all $r\in\mathbb{R}_{+}$, which proves that $W$ is Lipschitz continuous.
  2. (2)
    A drawback of example (1) is that $W_{K_{\rho,\rho^{\prime}}}$ is not available in closed form. For a more practical choice, set $C:=2(K_{\rho,\rho^{\prime}}\vee 1)$ and define the Lipschitz continuous function $W:\mathbb{R}_{+}\to\mathbb{R}_{+}$ as
    $W(r):=\frac{\ln(r+1)}{C},\quad\text{ for }r\in\mathbb{R}_{+}.$
    Obviously, $W$ is the inverse of the strictly increasing function $V:\mathbb{R}_{+}\to\mathbb{R}_{+}$ given by
    $V(r):=\exp(Cr)-1,\quad\text{ for }r\in\mathbb{R}_{+}.$
    Setting $M:=C/2=K_{\rho,\rho^{\prime}}\vee 1$ and recalling (4.28), we can compute
    $1+V_{K_{\rho,\rho^{\prime}}}(r)$ $\leq 1+r(1+r)\exp(Mr)=1+\sum_{k=0}^{\infty}\frac{M^{k}r^{k+1}}{k!}+\sum_{k=0}^{\infty}\frac{M^{k}r^{k+2}}{k!}$
    $=1+\sum_{k=1}^{\infty}\frac{M^{k-1}r^{k}}{k!}+\sum_{k=2}^{\infty}\frac{M^{k-2}r^{k}}{k!}=1+\frac{1}{M}\sum_{k=1}^{\infty}\frac{M^{k}r^{k}}{k!}+\frac{1}{M^{2}}\sum_{k=2}^{\infty}\frac{M^{k}r^{k}}{k!}$
    $\leq 1+\sum_{k=1}^{\infty}\frac{2^{k}M^{k}r^{k}}{k!}=\exp(2Mr)=\exp(Cr),$
    for all $r\in\mathbb{R}_{+}$. Therefore, we obtain $V_{K_{\rho,\rho^{\prime}}}(r)\leq V(r)$ for all $r\in\mathbb{R}_{+}$, which implies (4.51).
@@ 4.3. Monotonicity of term structures
where we recall that the state space is given by $H=H_{\rho}^{m+1}$. We say that the closed convex cone $K$ is invariant for the SPDE (4.25) if, for each $h_{0}\in K$, we have $\eta\in K$ up to an evanescent set, where $\eta$ denotes the mild solution to the SPDE (4.25) with $\eta_{0}=h_{0}$. Note that the cone $K$ is invariant for the shift semigroup; that is $\mathcal{S}_{t}K\subset K$ for all $t\geq 0$. The invariance of closed convex cones for SPDEs has been investigated in a general setup in [56, 57] and in [25] for the special case where $K$ is the cone of non-negative functions. The proof of the following theorem is analogous to that of [57, Corollary 8.6] and, therefore, omitted.
@@
$\quad-\int_{E}\big(\gamma^{i}(h,x)(\xi)-\gamma^{j}(h,x)(\xi)\big)F({\rm d}x)\geq 0,\quad\text{outside of a set of $(\mathbb{P}\otimes{\rm d}t)$-measure zero.}$
@@
for $(\mathbb{P}\otimes{\rm d}t)$-a.e. $(\omega,t)\in\Omega\times\mathbb{R}_{+}$. Recalling that $\psi$ is $(-1,+\infty)$-valued and $c^{i}$ is $1,+\infty)-$valued, conditions (4.58), (4.59), (4.61), (4.62) imply that (4.56) is fulfilled. Consequently, in view of Theorem 4.16, the closed convex cone $K$ is invariant for the HJMM SPDE (4.25). ∎
@@ 4.4. Existence of affine realizations
where $L$ is an $\mathbb{R}^{d}$-valued Lévy process, $A$ denotes the generator of a pseudo-contractive $C_{0}$-semigroup $(\mathcal{S}_{t})_{t\geq 0}$ on $H$ and $\alpha:\Omega\times\mathbb{R}_{+}\times H\to H$ and $\beta:\Omega\times\mathbb{R}_{+}\times H\to L_{2}(\mathbb{R}^{d},H)$. For each $j=1,\ldots,d$, we introduce the mapping $\beta^{j}:\Omega\times\mathbb{R}_{+}\times H\to H$ given by $\beta^{j}(\omega,t,h):=\beta(\omega,t,h)e_{j}$. We assume that conditions (4.13), (4.14) and (4.16), (4.17) from Theorem 4.2 are satisfied, which ensures existence and uniqueness of mild solutions to the Lévy-driven SPDE (4.64).
@@
Let $V\subset D(A)$ be an $n$-dimensional AA-invariant subspace such that the following conditions are fulfilled:
@@
$\phi^{x_{0}}(t):=\mathcal{S}_{t}x_{0}+\int_{0}^{t}\mathcal{S}_{t-s}\alpha_{1}(s){\rm d}s,\quad\text{ for all }t\in\mathbb{R}_{+}.$
@@
see for instance [50, Theorem 4.2.4]. In particular, it follows that $A\phi^{x_{0}}:\mathbb{R}_{+}\to H$ is continuous. Condition (4.67) is satisfied due to (4.71). Furthermore, for all $(\omega,t)\in\Omega\times\mathbb{R}_{+}$ and $v\in V$, we obtain
@@
Suppose that the SPDE (4.64) has an $n$-dimensional affine realization induced by $(\phi,T)$. Let $\ell\in L(H,\mathbb{R}^{n})$ be such that the linear operator $U:=\ell T\in L(\mathbb{R}^{n})$ is an isomorphism and let $R\in L(\mathbb{R}^{n},V)$ be the isomorphism $R:=TU^{-1}$. Then, for each $x_{0}\in D(A)$, the following hold:
@@
$Y_{t}^{x_{0}}=U^{-1}\ell(X_{t}^{x_{0}}-\phi^{x_{0}}(t)),\quad\text{ for all }t\in\mathbb{R}_{+}.$
@@
    (4.80) $\varphi^{h_{0},i}(t):=\mathcal{S}_{t}h_{0}^{i}-h_{0}^{i}(t)e^{\delta_{i}\cdot}+\frac{c_{i}^{2}}{2\delta_{i}^{2}}\big(e^{2\delta_{i}t}-1\big)\big(e^{\delta_{i}\cdot}-1\big)e^{\delta_{i}\cdot},\;\text{ for all }t\in\mathbb{R}_{+}\text{ and }i=0,1,\ldots,m;$
@@
$\phi^{h_{0}}(t)=\mathcal{S}_{t}h_{0}+\int_{0}^{t}\mathcal{S}_{t-s}\alpha_{1}\,{\rm d}s,\qquad\text{ for all }t\in\mathbb{R}_{+}.$
Note that $U:=\ell T\in L(\mathbb{R}^{m+1})$ is simply given by $U={\rm Id}$. The isomorphism $R:=TU^{-1}\in L(\mathbb{R}^{m+1},V)$ then reduces to $R=T$ and, therefore, $R\ell\in L(H,V)$ is given by
@@
$\int_{0}^{t}\mathcal{S}_{t-s}\alpha_{1}\,{\rm d}s=\bigg(\frac{c_{i}^{2}}{2\delta_{i}^{2}}\Big(\big(e^{2\delta_{i}t}-1\big)e^{\delta_{i}\cdot}-2\big(e^{\delta_{i}t}-1\big)\Big)e^{\delta_{i}\cdot}\bigg)_{i=0,1,\ldots,m}.$
@@
$R\ell\int_{0}^{t}\mathcal{S}_{t-s}\alpha_{1}\,{\rm d}s=\bigg(\frac{c_{i}^{2}}{2\delta_{i}^{2}}\big(e^{2\delta_{i}t}-2e^{\delta_{i}t}+1\big)e^{\delta_{i}\cdot}\bigg)_{i=0,1,\ldots,m}.$
@@
$\frac{{\rm d}}{{\rm d}\xi}\varphi^{h_{0}}(t)=\bigg(\mathcal{S}_{t}\frac{{\rm d}}{{\rm d}\xi}h_{0}^{i}-\delta_{i}h_{0}^{i}(t)e^{\delta_{i}\cdot}+\frac{c_{i}^{2}}{2\delta_{i}}\big(e^{2\delta_{i}t}-1\big)\big(2e^{2\delta_{i}\cdot}-e^{\delta_{i}\cdot}\big)\bigg)_{i=0,1,\ldots,m},$
@@
Let us define $\mathcal{X}^{\mathrm{fin}}_{1}:=\bigcup_{n\geq 1}\mathcal{X}^{n}_{1}$, so that $\mathcal{X}_{1}=\overline{\mathcal{X}^{\mathrm{fin}}_{1}}$, with the bar denoting the closure in Émery’s semimartingale topology. For each $n\in\mathbb{N}$, we denote by $\mathcal{X}^{\mathrm{fin}}_{1,[0,n]}:=\{X_{\cdot\wedge n}:X\in\mathcal{X}^{\mathrm{fin}}_{1}\}$ the set of all elements of $\mathcal{X}^{\mathrm{fin}}_{1}$ stopped at $n$ and we define $\mathcal{X}_{1,[0,n]}$ analogously. Since the set $\mathcal{X}^{\mathrm{fin}}_{1}$ is stable under stopping and stopping is a continuous operation in the semimartingale topology (see [47, Remark II.6.c]), it holds that $\overline{\mathcal{X}^{\mathrm{fin}}_{1,[0,n]}}=\mathcal{X}_{1,[0,n]}$. Moreover, it is easy to see that the sets $1+\mathcal{X}^{\mathrm{fin}}_{1,[0,n]}$ satisfy the requirements of [44, Definition 1.1], since fork-convexity can be shown analogously to [18, Lemma 2.1]. Suppose first that NUPBR holds. Then, for each $n\in\mathbb{N}$, the set $\mathcal{X}_{1}(n)$ is bounded in probability and therefore, by [44, Theorem 1.7], there exists $\widehat{X}^{n}\in\mathcal{X}_{1,[0,n]}$ with $\widehat{X}^{n}>-1$ such that $(1+X)/(1+\widehat{X}^{n})$ is a supermartingale, for every $X\in\mathcal{X}_{1,[0,n]}$. For each $n\in\mathbb{N}$, let us denote $D^{n}:=1/(1+\widehat{X}^{n})$. Similarly as in the proof of [14, Proposition 1], we show that the elements $D^{n}$, $n\in\mathbb{N}$, can be concatenated into a supermartingale deflator. For all $t\geq 0$, let $n(t):=\min\{n\in\mathbb{N}:n>t\}$ and define the càdlàg process
(A.1) $D_{t}:=\prod_{k=1}^{n(t)}\frac{D^{k}_{k\wedge t}}{D^{k}_{(k-1)\wedge t}},\qquad\text{ for all }t\geq 0.$
@@
Observe that this definition implies that, if $t\in(m-1,m]$ for some $m\in\mathbb{N}$, then $D_{t}=\prod_{k=1}^{m}(D^{k}_{k\wedge t}/D^{k}_{k-1})$. This follows from (A.1) by noting that if $t\in(m-1,m)$ then $n(t)=m$, while if $t=m$ then $n(t)=m+1$. Let $X\in\mathcal{X}_{1}$ and $s<t$. Suppose first that $t\in(s,n(s)]\subseteq(n(s)-1,n(s)]$. In this case, we can compute
@@
$\mathbb{E}\left[D_{t}(1+X_{t})|\mathcal{F}_{s}\right]$ $=\mathbb{E}\left[\prod_{k=1}^{n(s)}\frac{D^{k}_{k\wedge t}}{D^{k}_{k-1}}(1+X_{t})\Bigg|\mathcal{F}_{s}\right]$
$=\prod_{k=1}^{n(s)-1}\frac{D^{k}_{k}}{D^{k}_{k-1}}\,\mathbb{E}\left[\frac{D^{n(s)}_{t}}{D^{n(s)}_{n(s)-1}}(1+X_{t})\Bigg|\mathcal{F}_{s}\right]$
$\leq\prod_{k=1}^{n(s)-1}\frac{D^{k}_{k}}{D^{k}_{k-1}}\frac{D^{n(s)}_{s}(1+X_{s})}{{D^{n(s)}_{n(s)-1}}}=\prod_{k=1}^{n(s)}\frac{D^{k}_{k\wedge s}}{D^{k}_{k-1}}(1+X_{s})=D_{s}(1+X_{s}),$
@@
where the inequality follows from the fact that $D^{n(s)}(1+X)$ is a supermartingale on $[0,n(s)]$. This shows that the supermartingale property holds for $t\in(s,n(s)]$. To prove that $D$ is a supermartingale deflator, it suffices to show that the supermartingale property holds for $t\in(s,n(s)+l]$, for all $l\in\mathbb{N}$. We proceed by induction on $l$. Suppose that the supermartingale property holds for $t\in(s,n(s)+l]$, for some $l\in\mathbb{N}$. By the induction hypothesis, it suffices to prove the property for $t\in(n(s)+l,n(s)+l+1]$. To this effect, we compute
$\mathbb{E}\left[D_{t}(1+X_{t})|\mathcal{F}_{s}\right]$ $=\mathbb{E}\left[\mathbb{E}\left[D_{t}(1+X_{t})|\mathcal{F}_{n(s)+l}\right]\big|\mathcal{F}_{s}\right]$
$=\mathbb{E}\left[D_{n(s)+l}\,\mathbb{E}\left[\frac{D^{n(s)+l+1}_{t}}{D^{n(s)+l+1}_{n(s)+l}}(1+X_{t})\Bigg|\mathcal{F}_{n(s)+l}\right]\Bigg|\mathcal{F}_{s}\right]$
$\leq\mathbb{E}\left[D_{n(s)+l}(1+X_{n(s)+l})|\mathcal{F}_{s}\right]\leq D_{s}(1+X_{s}),$
@@
where the first inequality follows from the supermartingale property of $D^{n(s)+l+1}(1+X)$ on the interval $[0,n(s)+l+1]$ and the second inequality from the induction step. Since $D_{0}=1$, this completes the proof that the process $D$ defined in (A.1) is a supermartingale deflator in the sense of Definition 2.4. We now prove that $1/D\in 1+\mathcal{X}_{1}$. To this effect, note that
$\widehat{X}_{t}:=\frac{1}{D_{t}}-1=\prod_{k=1}^{n(t)}\frac{1+\widehat{X}^{k}_{k\wedge t}}{1+\widehat{X}^{k}_{(k-1)\wedge t}}-1,\qquad\text{ for all }t\geq 0.$